# Synergistic Dynamical Decoupling and Circuit Design for Enhanced Algorithm Performance on Near-Term Quantum Devices

### v1

@@ II.3 Synergistic design approach
This section describes a synergistic design approach for maximizing the performance and robustness of algorithms on near term quantum devices. This approach acknowledges the critical interplay between the hardware’s capabilities and the design choices made in the software implementation. The quality of algorithm implementation directly affects the performance. Key aspects include the efficiency of transpilation processes, specific gate types used, and the overall symmetry of the algorithm structure. For instance, studies have shown that the algorithm-oriented qubit mapping (AOQMAP) method [44] offers advantages in transpilation for variational quantum algorithms (VQAs) [45] compared to popular compilers such as Qiskit [46] and Tket [47] by introducing fewer two-qubit gates, maintaining a shallower circuit depth, and promoting higher symmetry [48].
@@
Figure 2: Two types of DD sequences: (a) CPMG and (b) XY4. The delay time $t$ represents the idle time of the qubit, minus the duration of the corresponding X or Y pulses.
@@
Figure 3: Two implementations of a three-qubit QAOA: (a) CX-implementation, and (b) CZ-implementation, both are decomposed and optimized using Qiskit’s transpiler with optimization level 3. Highlighted yellow boxes represent CPMG sequences.
In this study, we utilize the AOQMAP method [44] to efficiently map circuits onto hardware, aiming to minimize SWAP gates and circuit depth on linear topologies. Subsequently, we examine two implementations of QAOA for portfolio optimization on CX-based IBM QPUs. These implementations differ in their choice of gate decompositions within the algorithms. The first implementation, referred to as the CX implementation, directly decomposes the gates in QAOA into basis gates of the QPUs using Qiskit’s transpiler with optimization level 3. In comparison, the second implementation, referred to as the CZ implementation, initially decomposes gates in algorithms into CZ and single qubit gates. Then, Qiskit’s transpiler with optimization level 3 is used to perform optimization and decomposition into basis gates of QPUs. Previous studies have demonstrated that this CZ decomposition approach outperforms CX decomposition for ZZ and ZZ-SWAP gates in QAOA on IBM QPUs [43]. During implementation, we also explore different optimization level settings in Qiskit and investigate their impact. To ensure consistency in evaluations, we employ identical benchmark circuits and parameters. Additionally, we consistently use 30000 shots for each demonstration. Within the Qiskit framework, we default to using optimization level 3.
We also investigate the effectiveness of two well-established DD sequences: CPMG and XY4. As illustrated in Fig. 2, the CPMG sequence applies two X pulses separated by a delay of $\frac{t}{2}$, with additional delays of $\frac{t}{4}$ at the beginning and end. The parameter $t$ represents the time interval during which the qubit remains idle, excluding the duration of single qubit pulses, in the case of CPMG, two X pulses. In comparison, the XY4 sequence utilizes two X and two Y pulses, each separated by a delay of $\frac{t}{4}$, with additional delays of $\frac{t}{8}$ at the beginning and end. Additionally, the “alap” (as late as possible) scheduling method, which schedules the stop time of instructions as late as possible, is used for scheduling gates and inserting DD sequences throughout our study. Figure 3 showcases the resulting implementations of a three-qubit QAOA with CPMG sequence on a 27-qubit QPU ibm_kolkata using both CX and CZ implementations. Compared to CX implementation, CZ implementation employs all the same directed CX gates. Additionally, we observe an X gate inserted between control qubits of CX gates in the CZ implementation, potentially suppressing idle errors and improving performance. Moreover, CZ implementation shows improved symmetry than CX implementation. By simultaneously optimizing both hardware and software aspects through careful algorithm design, efficient transpilation techniques, and DD sequences, it is possible to fully exploit the capabilities of near term quantum algorithms.
@@ III.1 Impact of hardware factors
Our investigation begins by examining how hardware characteristics affect algorithm performance and DD effectiveness. The CPMG sequence is chosen for our study due to its widespread adoption and relative simplicity, allowing us to gain fundamental insights. We analyze these factors using extensive datasets. The experiments involve varying the qubit counts from 3 to 12 and investigating different combinations of algorithm implementations, including the CX and CZ versions of QAOA, as well as optimization levels 1 and 3 within the Qiskit framework. We conduct these experiments using eight QPUs, which consist of four 27-qubit devices: ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen, and four 127-qubit devices: ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke. These extensive datasets provide a solid foundation for analyzing the impact of hardware factors on algorithm performance.
@@ III.1.1 Circuit fidelity
Figure 4: Impact of circuit fidelity on algorithm performance and DD effectiveness. Higher values of $\mathrm{NAR}_{\mathrm{B}}$, $\mathrm{NAR}_{\mathrm{DD}}$, $\mathrm{NSP}_{\mathrm{B}}$, and $\mathrm{NSP}_{\mathrm{DD}}$ indicate better performance on actual quantum devices. Values exceeding unity indicate that the performance achieved on the IBM quantum hardware surpasses the results obtained from the noise-free simulation. Positive values of $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$ demonstrate improvements due to DD. The CPMG sequence is used for all data points. Each line in the graph represents a linear fit of the data. The reported $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$ are 85.55% and 66.8%, respectively.
@@
$\mathrm{NAR}_{\mathrm{B}}$ 0.610 $y=\phantom{-}0.572x+0.317$ $8090$
$\mathrm{NAR}_{\mathrm{DD}}$ 0.687 $y=\phantom{-}0.506x+0.430$ $8270$
$\mathrm{NSP}_{\mathrm{B}}$ 0.556 $y=\phantom{-}0.558x+0.270$ $5300$
$\mathrm{NSP}_{\mathrm{DD}}$ 0.631 $y=\phantom{-}0.537x+0.358$ $5210$
$\Delta_{\mathrm{NAR}}$ 0.077 $y=-0.065x+0.111$ $-0.209$ 0.00075
$\Delta_{\mathrm{NSP}}$ 0.075 $y=-0.021x+0.086$ $-0.027$ 0.66997
where $f_{G_{s}}$, $f_{G_{t}}$, and $f_{G_{m}}$ denote fidelities of a single qubit gate $G_{s}$, a two qubit gate $G_{t}$, and the measurement $G_{m}$, respectively, in the circuit. We focus on circuits with fidelities exceeding 0.01 and present results in Fig. 4, including $\mathrm{NAR}_{\mathrm{B}}$, $\mathrm{NAR}_{\mathrm{DD}}$, $\mathrm{NSP}_{\mathrm{B}}$, $\mathrm{NSP}_{\mathrm{DD}}$, $\Delta_{\mathrm{NAR}}$, and $\Delta_{\mathrm{NSP}}$, as defined in Sec. II.1.2. The measured data and the corresponding circuit fidelity are fitted using a linear function. The correlation coefficient $C_{r}$ and $p$-value are computed to assess the quality of linear approximation. $C_{r}$ quantifies the magnitude and direction of a linear relationship between two variables. It spans from $-1$ (indicating a perfect negative association) to 1 (indicating a perfect positive association), with 0 denoting no association. The absolute value of $C_{r}$ indicates the correlation strength, which can be interpreted according to established conventions: very strong correlation (|$\absolutevalue{C_{r}}$ | between 0.9 and 1.0), strong correlation (|$\absolutevalue{C_{r}}$ | between 0.7 and 0.9), moderate correlation (|$\absolutevalue{C_{r}}$ | between 0.4 and 0.7), weak correlation (|$\absolutevalue{C_{r}}$ | between 0.2 and 0.4), and very weak correlation (|$\absolutevalue{C_{r}}$ | between 0 and 0.2). However, it is important to note that correlation does not imply causation. The $p$-value is a complementary statistical measure that evaluates the strength of evidence against the null hypothesis of no correlation. A low $p$-value (typically below 0.05) suggests a statistically significant correlation, while a high $p$-value indicates that the correlation might be due to random chance. Linear fitting takes correlation a step further by determining the equation of a best fit line, enabling predictions based on the observed linear relationship.
Figure 4 depicts a general trend of improved algorithm performance with increasing circuit fidelity. DD sequences enhance both NAR and NSP on average. However, NSP exhibits a wider range of variation for a given circuit fidelity compared to approximation ratio, particularly at lower fidelities. Table 2 presents the average value, fitted function, correlation coefficient, and $p$-value for each metric. A stronger correlation is observed between NAR and circuit fidelity compared to NSP, suggesting a more pronounced dependence of NAR on fidelity. In contrast, the correlation between $\Delta_{\mathrm{NSP}}$ and circuit fidelity is very weak, as evidenced by the low value of $C_{r}$ and high $p$-value. This suggests that the observed decrease in DD effectiveness might be due to random fluctuations or other factors not captured by circuit fidelity alone. The average improvement in NAR and NSP due to DD sequences is approximately 0.08. Moreover, the negative coefficients of the fitted lines for $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$ suggest a potential decrease in DD effectiveness as circuit fidelity increases. Additionally, EMSR calculated with the approximation ratio ($\mathrm{EMSR}_{\mathrm{AR}}$) is 85.55%, while EMSR based on success probability ($\mathrm{EMSR}_{\mathrm{SP}}$) is 66.8%. On average, DD sequences demonstrate a success rate of 76.18% in improving algorithm performance, highlighting their robustness in mitigating errors in practical quantum algorithms.
@@ III.1.2 Schedule duration
Figure 5: Influence of logarithmic transformation of circuit schedule duration ($\ln(\tau)$) on algorithm performance and DD effectiveness with the same datasets as in Fig. 4.
@@
$\mathrm{NAR}_{\mathrm{B}}$ $y=-0.234x+1.124$ $-0.732$ 0
$\mathrm{NAR}_{\mathrm{DD}}$ $y=-0.194x+1.114$ $-0.696$ 0
$\mathrm{NSP}_{\mathrm{B}}$ $y=-0.247x+1.096$ $-0.519$ 0
$\mathrm{NSP}_{\mathrm{DD}}$ $y=-0.219x+1.111$ $-0.468$ 0
$\Delta_{\mathrm{NAR}}$ $y=\phantom{-}0.040x-0.010$ $2850$
@@
Figure 6: Influence of circuit fidelity and $\ln(\tau)$ on algorithm performance and DD effectiveness: (a) $\mathrm{NAR}_{\mathrm{DD}}$, (b) $\mathrm{NSP}_{\mathrm{DD}}$, (c) $\Delta_{\mathrm{NAR}}$, and (d) $\Delta_{\mathrm{NSP}}$.
We now investigate the influence of schedule duration $\tau$, and in particular the logarithmic transformation of schedule duration $\ln(\tau)$, on algorithm performance and DD effectiveness. Schedule duration reflects the total time required to execute a quantum circuit and depends on the number and execution time of individual gates. Shorter durations potentially improve circuit fidelity by reducing the system’s exposure to decoherence errors, but achieving them necessitates faster gates, which can be hardware limited.
Figure 5 depicts the impact of $\ln(\tau)$ on the defined metrics using the same datasets as in Fig. 4. We observe that the algorithm performance degrades with increasing schedule duration, while DD effectiveness improves. However, $\Delta_{\mathrm{NSP}}$ exhibits larger fluctuations for longer durations, suggesting that while DD sequences mitigate decoherence errors, potentially improving performance at longer durations, they could also introduce other error mechanisms such as operation errors that counteract this improvement. Table 3 summarizes corresponding parameters. The coefficients of linear function for $\mathrm{NAR}_{\mathrm{B}}$, $\mathrm{NAR}_{\mathrm{DD}}$, $\mathrm{NSP}_{\mathrm{B}}$, and $\mathrm{NSP}_{\mathrm{DD}}$ indicate a suppressed decay of performance with increasing schedule duration by applying DD sequences. The correlation coefficients between these metrics and $\ln(\tau)$ further support the effectiveness of DD sequences in reducing the dependence of performance (both NAR and NSP) on schedule duration. $\Delta_{\mathrm{NAR}}$ exhibits a statistically weak correlation with $\ln(\tau)$, while $\Delta_{\mathrm{NSP}}$ shows a very weak correlation. This observation aligns with the findings for circuit fidelity. However, $\Delta_{\mathrm{NSP}}$ appears more sensitive to schedule duration compared to circuit fidelity, as indicated by a larger absolute correlation coefficient (|$\absolutevalue{C_{r}}$ |) and lower $p$-value.
Figures 6(a-d) illustrate the impact of circuit fidelity and schedule duration on $\mathrm{NAR}_{\mathrm{DD}}$, $\mathrm{NSP}_{\mathrm{DD}}$, $\Delta_{\mathrm{NAR}}$, and $\Delta_{\mathrm{NSP}}$, respectively. As observed in Fig. 6(a), $\mathrm{NAR}_{\mathrm{DD}}$ exhibits degradation with increasing logarithmic schedule duration ($\ln(\tau)$) and decreasing circuit fidelity. High performance is concentrated in the region where the schedule duration $\tau$ is below $e^{2.1}$ dt and circuit fidelity surpasses 0.5. Conversely, low performance is primarily observed for $\tau$ exceeding $e^{2.1}$ dt and fidelities below 0.5. A similar trend is evident for $\mathrm{NSP}_{\mathrm{DD}}$ in Fig. 6(b). However, unlike $\mathrm{NAR}_{\mathrm{DD}}$, achieving a high $\mathrm{NSP}_{\mathrm{DD}}$ value remains feasible even for longer schedule durations and lower fidelities. This suggests that $\mathrm{NSP}_{\mathrm{DD}}$ is less sensitive to these factors compared to $\mathrm{NAR}_{\mathrm{DD}}$. Figures 6(c) and 6(d) further demonstrate the effectiveness of DD sequences, particularly at longer durations (higher $\ln(\tau)$). This is potentially due to the ability of DD sequences to mitigate decoherence errors that become more prominent at these timescales.

### v2

@@ II.3 Synergistic Design Approach
This section describes a synergistic design approach for maximizing the performance and robustness of algorithms on near-term quantum devices. This approach acknowledges the critical interplay between the hardware’s capabilities and the design choices made in the software implementation. The quality of algorithm implementation directly affects the performance. Key aspects include the efficiency of transpilation processes, specific gate types used, and the overall symmetry of the algorithm structure. For instance, studies have shown that the algorithm-oriented qubit mapping (AOQMAP) method [42] offers advantages in transpilation for variational quantum algorithms (VQAs) [49] compared to popular compilers such as Qiskit [50] and Tket [51] by introducing fewer two-qubit gates, maintaining a shallower circuit depth, and promoting higher symmetry [43].
In this study, we utilize the AOQMAP method [42] to efficiently map circuits onto hardware, aiming to minimize SWAP gates and circuit depth on linear topologies. Subsequently, we examine two implementations of QAOAs for portfolio optimization on CX-based IBM QPUs. These implementations differ in their choice of gate decompositions within the algorithms. The first implementation, referred to as CX implementation, directly decomposes the gates in the QAOA into basis gates of the QPUs using Qiskit’s transpiler with optimization level 3. In comparison, the second implementation, referred to as CZ implementation, initially decomposes gates in algorithms into CZ and single-qubit gates. Then, Qiskit’s transpiler with optimization level 3 is used to perform optimization and decomposition into basis gates of QPUs. Previous studies have demonstrated that this CZ decomposition approach outperforms CX decomposition for ZZ and ZZ-SWAP gates in the QAOA on IBM QPUs [44]. During implementation, we also explore different optimization level settings in Qiskit and investigate their impact. To ensure consistency in evaluations, we employ identical benchmark circuits and parameters. Additionally, we consistently use 30,000 shots for each demonstration. Within the Qiskit framework, we default to using optimization level 3.
We also investigate the effectiveness of two well-established DD sequences: CPMG and XY4. As illustrated in Fig. 2, the CPMG sequence applies two X pulses separated by a delay of $\frac{t}{2}$, with additional delays of $\frac{t}{4}$ at the beginning and end. The parameter $t$ represents the time interval during which the qubit remains idle, excluding the duration of single-qubit pulses and, in the case of CPMG, two X pulses. In comparison, the XY4 sequence utilizes two X and two Y pulses, each separated by a delay of $\frac{t}{4}$, with additional delays of $\frac{t}{8}$ at the beginning and end. Additionally, the “alap” (as late as possible) scheduling method, which schedules the stop time of instructions as late as possible, is used for scheduling gates and inserting DD sequences throughout our study. Figure 3 showcases the resulting implementations of a three-qubit QAOA with a CPMG sequence on a 27-qubit QPU ibm_kolkata using both CX and CZ implementations. Compared to CX implementation, CZ implementation employs all the same directed CX gates. Additionally, we observe an X gate inserted between control qubits of CX gates in the CZ implementation, potentially suppressing idle errors and improving performance. Moreover, CZ implementation shows improved symmetry compared to CX implementation. By simultaneously optimizing both hardware and software aspects through careful algorithm design, efficient transpilation techniques, and DD sequences, it is possible to fully exploit the capabilities of near-term quantum algorithms.
@@
Figure 2: Two types of dynamical decoupling (DD) sequences: (a) Carr–Purcell–Meiboom–Gill (CPMG) and (b) XY4. The delay time $t$ represents the idle time of the qubit minus the duration of the corresponding X or Y pulses.
@@
Figure 3: Two implementations of a three-qubit QAOA: (a) CX implementation and (b) CZ implementation, where both are decomposed and optimized using Qiskit’s transpiler with optimization level 3. Highlighted yellow boxes represent CPMG sequences.
@@ III.1 Impact of Hardware Factors
Our investigation begins by examining how hardware characteristics affect algorithm performance and DD effectiveness. The CPMG sequence is chosen for our study due to its widespread adoption and relative simplicity, allowing us to gain fundamental insights. We analyze these factors using extensive datasets. The experiments involve varying the qubit counts from 3 to 12 and investigating different combinations of algorithm implementations, including the CX and CZ versions of the QAOA, as well as optimization levels 1 and 3 within the Qiskit framework. We conduct these experiments using eight QPUs, which consist of four 27-qubit devices—ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen—and four 127-qubit devices: ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke. These extensive datasets provide a solid foundation for analyzing the impact of hardware factors on algorithm performance.
@@ III.1.1 Circuit Fidelity
where $f_{G_{s}}$, $f_{G_{t}}$, and $f_{G_{m}}$ denote fidelities of a single-qubit gate $G_{s}$, a two-qubit gate $G_{t}$, and the measurement $G_{m}$, respectively, in the circuit. We focus on circuits with fidelities between 0.01 and 1. This broader range of fidelities establishes a solid basis for investigating the effectiveness of DD sequences under realistic noise conditions. We analyze various metrics defined in Section II.1.2, including $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$, $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$. The measured data and the corresponding circuit fidelity are fitted using a linear function. The correlation coefficient $C_{r}$ and $p$-value are computed to assess the quality of linear approximation. $C_{r}$ measures the strength and direction of the linear relationship, ranging from $-1$ (perfect negative correlation) to 1 (perfect positive correlation), with 0 indicating no association. The absolute value of $C_{r}$ reflects the correlation strength: very strong (0.9–1.0), strong (0.7–0.9), moderate (0.4–0.7), weak (0.2–0.4), and very weak (0–0.2). It is important to note that correlation does not imply causation. The $p$-value is a complementary statistical measure that evaluates the strength of evidence against the null hypothesis of no correlation. A low $p$-value (typically below 0.05) suggests a statistically significant correlation, possibly not due to random chance. Linear fitting builds on correlation by determining the best-fit line equation, enabling predictions based on the observed relationship.
@@
Figure 4: Impact of circuit fidelity on algorithm performance and DD effectiveness: (a) normalized approximation ratio (NAR), (b) normalized success probability (NSP), (c) improvement in NAR after applying DD ($\Delta_{\mathrm{\scriptscriptstyle NAR}}$), and (d) improvement in NSP after applying DD ($\Delta_{\mathrm{\scriptscriptstyle NSP}}$). Higher values of $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$, and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ indicate better performance on actual quantum devices. Values exceeding unity indicate that the performance achieved on the IBM quantum hardware surpasses the results obtained from the noise-free simulation. Positive values of $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ demonstrate improvements due to DD. The CPMG sequence is used for all data points. Each line in the graph represents a linear fit of the data. The reported $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$ are 85.55% and 66.8%, respectively.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$ 0.610 $y=\phantom{-}0.572x+0.317$ $8090$
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$ 0.687 $y=\phantom{-}0.506x+0.430$ $8270$
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$ 0.556 $y=\phantom{-}0.558x+0.270$ $5300$
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ 0.631 $y=\phantom{-}0.537x+0.358$ $5210$
$\Delta_{\mathrm{\scriptscriptstyle NAR}}$ 0.077 $y=-0.065x+0.111$ $-0.209$ 0.00075
$\Delta_{\mathrm{\scriptscriptstyle NSP}}$ 0.075 $y=-0.021x+0.086$ $-0.027$ 0.66997
Figure 4 depicts a general trend of improved algorithm performance with increasing circuit fidelity. DD sequences enhance both the NAR and NSP on average. However, the NSP exhibits a wider range of variation for a given circuit fidelity compared to the approximation ratio, particularly at lower fidelities. Table 2 presents the average value, fitted function, correlation coefficient, and $p$-value for each metric. A stronger correlation is observed between NAR and circuit fidelity compared to NSP, suggesting a more pronounced dependence of NAR on fidelity. In contrast, the correlation between $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ and circuit fidelity is very weak, as evidenced by the low value of $C_{r}$ and high $p$-value. This suggests that the observed decrease in DD effectiveness might be due to random fluctuations or other factors not captured by circuit fidelity alone. The average improvement in NAR and NSP due to DD sequences is approximately 0.08. Moreover, the negative coefficients of the fitted lines for $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ suggest a potential decrease in DD effectiveness as circuit fidelity increases.
@@ III.1.2 Schedule Duration
We now investigate the influence of schedule duration $\tau$ and, in particular, the logarithmic transformation of schedule duration $\ln(\tau/\mathrm{dt})$ on algorithm performance and DD effectiveness. Schedule duration reflects the total time required to execute a quantum circuit and depends on the number and execution time of individual gates. Shorter durations potentially improve circuit fidelity by reducing the system’s exposure to decoherence errors, but achieving them necessitates faster gates, which can be hardware-limited.
@@
Figure 5: Influence of logarithmic transformation of circuit schedule duration ($\ln(\tau/\mathrm{dt})$) on algorithm performance and DD effectiveness with the same datasets as in Fig. 4: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, and (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$ $y={-0.238x+3.136}$ ${-0.731}$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$ $y={-0.196x+2.764}$ ${-0.691}$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$ $y={-0.254x+3.253}$ ${-0.524}$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ $y={-0.221x+2.975}$ ${-0.464}$ 0
@@
[…]
Figure 5 depicts the impact of $\ln(\tau/\mathrm{dt})$ on the defined metrics using the same datasets as in Fig. 4. We observe that the algorithm performance degrades with increasing schedule duration, while DD effectiveness improves. However, $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ exhibits larger fluctuations for longer durations, suggesting that while DD sequences mitigate decoherence errors, potentially improving performance at longer durations, they could also introduce other error mechanisms, such as operation errors, that counteract this improvement. Table 3 summarizes corresponding parameters. The coefficients of the linear function for $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$, and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ indicate a suppressed decay in performance with increasing schedule duration by applying DD sequences. The correlation coefficients between these metrics and $\ln(\tau/\mathrm{dt})$ further support the effectiveness of DD sequences in reducing the dependence of performance (both NAR and NSP) on schedule duration. $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ exhibits a statistically weak correlation with $\ln(\tau/\mathrm{dt})$, while $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ shows a very weak correlation. This observation aligns with the findings for circuit fidelity. However, $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ appears more sensitive to schedule duration compared to circuit fidelity, as indicated by a larger absolute correlation coefficient (|$\absolutevalue{C_{r}}$ |) and lower $p$-value.
Figures 6(a–d) illustrate the impact of circuit fidelity and schedule duration on $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$, $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, and $\Delta_{\mathrm{\scriptscriptstyle NSP}}$, respectively. As observed in Fig. 6(a), $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$ exhibits degradation with increasing logarithmic schedule duration ($\ln(\tau/\mathrm{dt})$) and decreasing circuit fidelity. High performance is concentrated in the region where the schedule duration $\tau$ is below $e^{10.5}\mathrm{dt}$ and circuit fidelity surpasses 0.5. Conversely, low performance is primarily observed for $\tau$ exceeding $e^{10.5}\mathrm{dt}$ and fidelities below 0.5. A similar trend is evident for $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ in Fig. 6(b). However, unlike $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, achieving a high $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ value remains feasible even for longer schedule durations and lower fidelities. This suggests that $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$ is less sensitive to these factors compared to $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$. Figures 6(c–d) further demonstrate the effectiveness of DD sequences, particularly at longer durations. This is potentially due to the ability of DD sequences to mitigate decoherence errors that become more prominent at these timescales.
@@
Figure 6: Influence of circuit fidelity and $\ln(\tau/\mathrm{dt})$ on algorithm performance and DD effectiveness: (a) $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$, (b) $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, and (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$.